OpenStudy (anonymous):

A water tank that holds 60 L of water can be emptied in 24 min. How long will it take to empty a water tank that holds 280 L of water?

hero (hero):

Set up proportion

OpenStudy (anonymous):

i did and get it wrong. let me try again

OpenStudy (anonymous):

got*

hero (hero):

\[\frac{60 L}{280 L} = \frac{24 \min}{x \min}\]

OpenStudy (anonymous):

ok i got it right thank you i got 112

hero (hero):

Cancel L/L and min/min to get: \[\frac{60}{280} = \frac{24}{x}\]

hero (hero):

Factors of one can be cancelled